OpenStudy (anonymous):

triangle XYZ where X(6, 3), Y(-2, 4), and Z(10, -2) becomes triangle X’Y’Z’ where X’(3, 1.5), Y’(-1, 2), and Z’(5, -1). What is the scale factor for this dilation?

OpenStudy (anonymous):

X'=mX \[\left(\begin{matrix}3 \\ 1.5\end{matrix}\right) = m \left(\begin{matrix}6 \\ 3\end{matrix}\right)\] \[\left(\begin{matrix}3 \\ 1.5\end{matrix}\right) = \left(\begin{matrix}6m \\ 3m\end{matrix}\right)\] 3=6m m=1/2 do you need further help?